# Visual Sparse Steering: Improving Zero-shot Image Classification with Sparsity Guided Steering Vectors

### v1

@@
Abstract
Steering vision foundation models at inference time without retraining or access to large labeled datasets is a desirable yet challenging objective, particularly in dynamic or resource-constrained settings. In this paper, we introduce Visual Sparse Steering (VS2), a lightweight, test-time method that guides vision models using steering vectors derived from sparse features learned by top-$k$ Sparse Autoencoders without requiring contrastive data. Specifically, VS2 surpasses zero-shot CLIP by 4.12% on CIFAR-100, 1.08% on CUB-200, and 1.84% on Tiny-ImageNet. We further propose VS2++, a retrieval-augmented variant that selectively amplifies relevant sparse features using pseudo-labeled neighbors at inference time. With oracle positive/negative sets, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Interestingly, VS2 and VS2++ raise per-class accuracy by up to 25% and 38%, respectively, showing that sparse steering benefits specific classes by disambiguating visually or taxonomically proximate categories rather than providing a uniform boost. Finally, to better align the sparse features learned through the SAE reconstruction task with those relevant for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS). By incorporating a prototype-alignment loss during SAE training, using labels only during training while remaining fully test-time unsupervised, PASS consistently, though modestly, outperforms VS2, achieving a 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32.
@@ 1 Introduction
Foundation Models (FMs) have exhibited strong generalization capabilities across a wide range of domains, from classification tasks to open-ended generation  clip ; sam ; llama2 ; llava ; gpt4 . Despite their impressive performance, these models largely function as black boxes, restricting their controllability, and reliability in practice. The standard approach to edit a model’s behavior involves first collecting a labeled dataset resembling the target objective, followed by a supervised fine-tuning (SFT) procedure upon the holistic zhang2023instruction ; dong2023abilities or partial (e.g., PEFT lester2021power ; lora ; ia3 ) parameter space. Nevertheless, this procedure is cumbersome and can become prohibitively expensive given the model scale. Consequently, Steering Vector (SV) methods turner2023activation ; zou2023representation ; liu2024incontext , as an alternative, have gained popularity in the Large Language Model (LLM) domain as they can effectively modulate the model’s generation at inference-time in a training-free manner. While SVs have shown promise in LLMs, their functionality in the vision domain remains unclear.
In this work, we investigate the functionality of steering vectors in vision foundation models (VFMs). In particular, we use zero-shot classification tasks as a testbed to explore how SVs can be leveraged to upweight relevant visual features. Conventional SVs operate by first constructing a directional vector within the latent space to indicate a desired change of the functionality. It then linearly combines the directional vector with the original latent embeddings to craft output embeddings. In LLMs, SVs rely on two symbolic anchors that are inherently interpretable to determine the functional direction of the SV. For example, a happiness SV can be extracted using the latent embedding arithmetic corresponding to Happy - Sad. The formulation of SVs in LLM poses two critical challenges in vision domain. Given an image at inference-time: (1) Unlike symbolic language tokens, there exists tremendous redundancy in visual representations; consequently, it is unclear what features should be steered; (2) Without inherently interpretable anchors (e.g., happy, sad), towards which direction to steer is not obvious.
@@
Figure 1: Overview of VS2 and VS2++. At inference, VS2 finds sparse latent concepts, and steers the original embedding towards the direction of amplifying those sparse features. When additional caching data is available, VS2 selectively enhances certain features while suppressing others in a contrastive manner.
To address the aforementioned challenges, we propose Visual Sparse Steering (VS2). Drawing inspiration from the interpretability provided by sparse auto-encoders (SAEs) hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , we hypothesize that the sparse features in a SAE’s latent space can capture key aspects of a visual embedding. VS2 upweights these features equally to construct a steering vector for classification. This equal upweighting of features is imprecise in that we expect different features to be more relevant than others for classification tasks. Despite this, we show that the simple equal upweighting of features still improves zero-shot classification.
To construct a more precise steering vector, we propose VS2++, which leverages additional unlabeled data to construct steering vectors for classification. For a given target image, VS2++ retrieves the Top-N most similar visual embeddings, and constructs positive and negative groups according to generated pseudo labels [^1] In practice, pseudo labels can be gathered using either unsupervised clustering methods or zero-shot classification results. The VS2++ SV is then constructed via subtracting the average SAE features of the negative group from the positive group. Intuitively, VS2++ upweights the SAE features most relevant for distinguishing the positive and negative groups. In summary, VS2 and VS2++ aim to enhance downstream classification performance while retaining the interpretability of SAEs. To further align the sparse features learned through the SAE reconstruction task with those important for downstream performance, we propose Prototype-Aligned Sparse Steering (PASS).
VS2 consistently outperforms all baseline methods across all datasets and ViT backbones, surpassing zero-shot CLIP  radford2021learning by 3.45% and 4.12% on CIFAR-100, 0.93% and 1.08% on CUB-200, and 1.50% and 1.84% on Tiny-ImageNet with ViT-B/32 and ViT-B/16 backbones, respectively, without any external data or supervision. When using external unlabeled image data, VS2++ achieves absolute top-1 gains over CLIP zero-shot of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet. Finally, by adding a lightweight prototype-alignment loss during SAE training (labels used only in training, still test-time unsupervised) PASS consistently, though slightly, outperforms VS2, offering a marginal 6.12% gain over VS2 only on CIFAR-100 with ViT-B/32. Our main contributions are:
  • We investigate the functionality of steering vectors in vision domain, and propose VS2 that leverages sparse feature learning to construct steering vectors for zero-shot classification.
  • We further propose VS2++ for scenarios where additional unlabeled data is available to selectively amplify important concepts. In this case, we use VS2 on pseudo-positive and -negative data to upweight relevant features for classification.
  • We propose Prototype-Aligned Sparse Steering (PASS), a lightweight method that introduces a prototype-alignment loss during SAE training using class labels only during training while remaining fully test-time unsupervised.
  • We empirically demonstrate that VS2, VS2++, and PASS can leverage SAEs to substantially improve zero-shot classification performance, outperforming baseline methods by a large margin under their corresponding regimes, while also retaining interpretability.
@@ 2 Related Work
Several traditional approaches exist for interpretability in vision models, including feature visualizationsimonyan2014visualising ; zeiler2014visualizing ; olah2017feature and network dissectionbau2017network ; oikarinen2022clip . Mechanistic interpretability seeks to systematically analyze and understand neural networks elhage2021mathematical ; olah2020zoom , but it faces challenges due to polysemantic neurons i.e. units that activate in response to multiple, seemingly unrelated inputs elhage2022toy . This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations elhage2022toy . Sparse Autoencoders (SAEs) have been shown to mitigate superposition by applying sparse dictionary learning to model internals sharkey2022taking ; bricken2023monosemanticity . Recent efforts have leveraged SAEs to uncover interpretable features within LLMs, revealing latent units tied to grammar rules, style patterns, and factual knowledge templeton2024scaling ; cunningham2023sparse ; gao2024scaling .  joshi2025identifiable refines this by training autoencoders on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision. While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
While some studies investigate SAEs for exploratory feature analysis or generative modeling in the visual domain hugofry2024multimodal ; Daujotas2024 ; bhalla2024interpreting ; stevens2025sparse , they primarily address concept discovery rather than performance enhancement. Other attempts in the vision domain try to create new SAE architectures for concept extraction  fel2025archetypal ; thasarathan2025universal . In contrast, our method leverages SAEs not only to interpret but also to actively steer vision models, yielding classification improvements alongside interpretable latent representations.
Similar to our line of work, joseph2025steering validate CLIP’s steerability and its efficacy in suppressing typographic attacks. Patch-SAElim2024sparse improves zero-shot classification by applying latent masking based on class-level activations. Specifically, it aggregates SAE activations from the training split to identify the top-$k$ most frequently activated latents per class, which are then used to mask active latents at test time. Our approach differs in a key aspect: we do not rely on activation aggregations from the same class in training but our method is instead fully test-time unsupervised. This design avoids reliance on class-conditioned statistics, making our approach more flexible and applicable to a wide range of tasks without requiring dataset-specific activation distributions. Specifically, our method (i) operates entirely in the vision latent space, (ii) requires no class labels, external vocabularies, or gradient updates, and (iii) improves zero-shot classification by amplifying sparse SAE features rather than replacing or suppressing activations.
@@
Steering vector (SV) methods turner2023activation ; park2023the ; hernandez2024linearity ; NIPS2013_9aa42b31 , also known as representation engineering zou2023representation , construct a directional vector (or task vector) and apply it in the latent space to change the target model’s behavior at inference-time. SVs have gained popularity in LLM/MLLM domain as it can improve generation security liu2024incontext , promoting truthfulness NEURIPS2023_81b83900 , mitigating hallucination vista , as well as improving inference efficiency liu2024incontext ; i2cl . Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts gavrikov2025can . In contrast, we differentiate from this line of work by focusing on Vision Foundation Models directly steering representations in the latent space without any external language signal.
Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions bayat2025steering ; makelov2024sparse . Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled concept pairs, making it more adaptable to general visual representations. Furthermore, its applicability to vision foundation models (VFM) remains unclear. The proposed visual sparse steering (VS2) method demonstrates that SV methods, combined with sparse autoencoders, can be utilized to enhance image features, improving zero-shot classification.
@@ 3 Method
3.1 Preliminaries
Sparse Autoencoder (SAE). A SAE consists of an encoder-decoder pair that maps an input vector ${\bm{x}}$ into a latent representation ${\bm{z}}$, then reconstructs ${\bm{x}}$ from ${\bm{z}}$. Formally, we define the encoder as:
${\bm{z}}=\mathrm{ReLU}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}\Bigr{)},$ (1)
where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input. The decoder then reconstructs ${\bm{x}}$ as:
$\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}},$ (2)
where ${\bm{W}}_{\mathrm{dec}}\in{\mathbb{R}}^{d\times n}$ is the decoder weight matrix. To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective is:
$\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ (3)
In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  gao2024scaling explicitly enforces the sparsity by using a top-$k$ selection mechanism:
${\bm{z}}=\mathrm{TopK}\Bigl{(}{\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})\Bigr{)},$ (4)
where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest. This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
@@ 3.2 Visual Sparse Steering
In this section, we introduce Visual Sparse Steering (VS2), a novel framework that leverages sparse features to steer neural representations toward meaningful semantic directions. Our approach targets two interconnected objectives, i.e., (1) explicitly identifying semantically meaningful concepts embedded within latent representations and (2) utilizing these interpretable concepts to define steering directions that enhance downstream task performance by emphasizing relevant semantic information and reducing spurious correlations. Unlike standard steering methods that rely on explicitly curated positive and negative examples, our approach utilizes Sparse Autoencoders (SAEs) to automatically discover and isolate sparse concepts that assist in defining a steering direction. By selectively amplifying these concepts within the latent space, VS2 generates conceptually grounded steering vectors that improve model interpretability, enhance classification accuracy, and robustly mitigate reliance on superficial or irrelevant cues.

### v2

# Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors

@@
Visual Sparse Steering (VS2): Unsupervised Adaptation for Image Classification using Sparsity-Guided Steering Vectors
Gerasimos Chatzoudis    Zhuowei Li    Gemma E. Moran    Hao Wang    Dimitris N. Metaxas
Machine Learning, ICML
Steering vision foundation models at test time, without updating foundation-model weights or using labeled target data, is a desirable yet challenging goal. We present Visual Sparse Steering (VS2), a lightweight, label-free adaptation method that constructs a steering vector from sparse features extracted by a Sparse Autoencoder (SAE) trained on unlabeled in-domain training-split activations of the vision encoder. VS2 offers three key advantages over existing test-time adaptation methods: (1) a feature-level intervention space in sparse SAE representations; (2) efficiency, requiring only a forward pass with no test-time optimization or backpropagation; and (3) a reliability diagnostic based on SAE reconstruction loss that can skip steering when reconstruction is poor, enabling safe fallback to the baseline, a capability not standard in conventional steering vectors and test-time adaptation methods. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones (ViT-B/32, ViT-B/16), VS2 improves zero-shot top-1 accuracy by 3.45–4.12%, 0.93–1.08%, and 1.50–1.84%, respectively, while remaining forward-only and adding minimal compute overhead. A retrieval-based upper-bound analysis suggests substantial headroom if task-relevant sparse features can be selected reliably, motivating future work on selective feature amplification for interpretable, efficient test-time steering.
@@ 1 Introduction
Adapting vision foundation models at test time, without labeled data or weight updates, is desirable but challenging. Many test-time adaptation (TTA) methods rely on backpropagation through large encoders, incurring substantial compute overhead  (Shu et al., 2022; Feng et al., 2023; Yoon et al., 2024). Steering vectors offer a lightweight alternative but typically require carefully curated contrastive examples, limiting practicality in vision  (Turner et al., 2023; Zou et al., 2023; Liu et al., 2024). This motivates two challenges: (i) Can we construct effective steering directions without contrastive supervision or test-time optimization? and (ii) Can we find representations that mitigate the redundancy and entanglement inherent in visual features?
Drawing inspiration from mechanistic interpretability work on sparse autoencoders (SAEs) (Elhage et al., 2022; Bricken et al., 2023), we hypothesize that sparse SAE latents capture salient aspects of the input and can provide a less redundant basis for intervention in visual representations. SAEs have been used primarily as an interpretability tool to extract human-interpretable features in trained foundation models. However, their potential as a control mechanism remains underexplored: it is unclear whether sparse features can be manipulated in a principled, label-free manner to improve downstream performance. This gap motivates our approach: rather than using SAE features for post-hoc analysis, we leverage them as an actionable intervention space for efficient, label-free test-time steering.
To this end, we propose Visual Sparse Steering (VS2): we train a top-$k$ SAE on unlabeled in-domain activations and, at test time, construct a steering vector by amplifying the image’s active sparse features and projecting the induced change back to embedding space. VS2 is forward-only, requiring a single SAE forward pass with no test-time optimization, backpropagation, or contrastive anchor examples. Across CIFAR-100, CUB-200, and Tiny-ImageNet with two CLIP backbones (ViT-B/32, ViT-B/16)  (Radford et al., 2021b), VS2 improves zero-shot top-1 accuracy by 3.45%-4.12%, 0.93%-1.08%, and 1.50%-1.84%, respectively, while adding $0.1\%$ compute overhead.
A practical challenge is distribution shift: if a test input lies outside the SAE’s training distribution, reconstruction quality can degrade, and the resulting sparse features may be unreliable, causing steering to underperform. To mitigate this, VS2 includes a built-in reliability diagnostic based on SAE reconstruction error. When reconstruction error is high, we skip steering and fall back to the original zero-shot CLIP prediction, avoiding harmful adaptation from misaligned sparse features. This provides a simple, model-internal mechanism for deciding when to apply steering, which is not typically available in standard steering-vector or test-time adaptation methods.
A natural question is whether all active sparse features should be amplified uniformly, or whether selectively amplifying task-relevant features could yield larger gains. To quantify this headroom, we introduce VS2++ as a retrieval-based upper-bound analysis: given an external unlabeled corpus, it retrieves nearest neighbors, forms positive and negative groups using CLIP pseudo-labels, and constructs a contrastive steering direction in SAE feature space. With oracle positives/negatives, VS2++ yields gains of up to 21.44% on CIFAR-100, 7.08% on CUB-200, and 20.47% on Tiny-ImageNet, illustrating substantial headroom when feature selection is accurate and motivating future work on robust selection under noisy pseudo-labels.
In summary, our contributions are:
  • Unsupervised sparse steering for vision. We introduce VS2, a lightweight, label-free method that trains a top-$k$ SAE on unlabeled in-domain activations and performs forward-only test-time steering without updating CLIP weights or using contrastive anchor examples.
  • Consistent gains with minimal overhead. Across CIFAR-100, CUB-200, and Tiny-ImageNet and two CLIP backbones, VS2 improves zero-shot top-1 accuracy while adding $0.1\%$s compute overhead and requiring no test-time optimization.
  • Reliability-aware fallback. We propose a reconstruction-error-based diagnostic that can conservatively disable steering and fall back to the baseline when the SAE representation is unreliable, reducing the risk of harmful interventions under distribution shift, at the cost of occasionally forgoing potential gains under a strict reconstruction-error threshold.
  • Headroom via selective amplification. We present VS2++ as a retrieval-based upper-bound analysis showing substantial potential gains when task-relevant sparse features can be selected accurately, motivating future work on robust feature selection under noisy pseudo-labels.
@@
Figure 1: Overview of VS2 and VS2++: at inference time, VS2 constructs a steering vector by equally amplifying sparse features, while VS2++ selectively enhances them when an external visual corpus is available.
@@ 2 Related Work
Several traditional approaches exist for interpretability in vision models, including feature visualization (Simonyan et al., 2014; Zeiler and Fergus, 2014; Olah et al., 2017) and network dissection (Bau et al., 2017; Oikarinen and Weng, 2022). Mechanistic interpretability seeks to systematically analyze and understand neural networks (Elhage et al., 2021; Olah et al., 2020), but it faces challenges due to polysemantic neurons i.e. units that activate in response to multiple, seemingly unrelated inputs (Elhage et al., 2022). This phenomenon arises from superposition, where networks encode more features than the available dimensions allow, forcing different concepts to share the same activations (Elhage et al., 2022). Sparse Autoencoders (SAEs) have been explored to mitigate superposition by applying sparse dictionary learning to model internals (Bricken et al., 2023). Recent efforts have leveraged SAEs to uncover interpretable features within LLMs  (Templeton, 2024; Cunningham et al., 2023; Gao et al., 2024a).  Joshi et al. (2025) train SAEs on embedding differences to disentangle multiple concept shifts, enabling precise interventions in model activations without requiring direct supervision. While these methods focus on language models, our work extends sparse steering to the vision domain, where applications of SAEs to vision models remain comparatively underexplored.
While SAEs have been explored for feature analysis, generative modeling, and concept disentanglement in the visual domain (Bhalla et al., 2024; Stevens et al., 2025; Surkov et al., 2024; Fel et al., 2025; Thasarathan et al., 2025), these works focus primarily on interpretability rather than downstream performance. In contrast, our method leverages SAEs to actively steer vision models in a label-free, test-time setting to improve classification. Related efforts like (Joseph et al., 2025) study CLIP’s steerability on typographic attacks, while Patch-SAE (Lim et al., 2024) improves classification accuracy via class-conditioned latent masking. Our approach, on the other hand, requires no class labels and avoids class-based activation aggregation during training, enabling broader applicability without reliance on external supervision or gradient updates.
@@
Steering Vector (SV) methods (Turner et al., 2023; Park et al., 2023; Hernandez et al., 2024; Mikolov et al., 2013), also known as representation engineering (Zou et al., 2023), construct a directional task vector and apply it in the latent space to change the target model’s behavior at inference time. In LLMs/MLLMs, SVs are used to enhance security (Liu et al., 2024), truthfulness (Li et al., 2023), reduce hallucinations (Li et al., 2025a), and improve efficiency (Li et al., 2025b). Interestingly, prior work has shown that in VLMs, visual cues are influenced by language, and that biases in the model’s response can be steered through simple natural language prompts (Gavrikov et al., 2025). In contrast, we focus on steering latent representations in vision models without any language input. Recent work has demonstrated that sparse representations can improve interpretability and disentanglement in steering directions (Bayat et al., 2025; Makelov, 2024). Unlike these methods, which rely on supervised contrastive examples or training data, our approach discovers meaningful sparse directions in vision models without requiring labeled positive/negative concept pairs, making it more adaptable to general visual representations.
Test-Time Adaptation.
Recent advances in improving the generalization of vision-language models have introduced a variety of prompt-tuning approaches, such as CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), and MaPLe (Khattak et al., 2023), along with adapter-based techniques like Tip-Adapter (Zhang et al., 2021) and CLIP-Adapter (Gao et al., 2024b). Despite their effectiveness, these methods typically assume access to a small number of labeled target-domain examples (often in a few-shot setting). In contrast, Test-Time Adaptation methods aim to improve robustness under distribution shift by adapting a pre-trained model using only unlabeled test samples at inference time (Liu et al., 2021; Sun et al., 2020; Wang et al., 2020; Gao et al., 2022). Specifically, Test-Time Prompt Tuning (TPT) (Shu et al., 2022) introduced entropy-minimizing prompt optimization over augmented views, inspiring subsequent methods such as DiffTPT (Feng et al., 2023), which leverages diffusion-based augmentations, and C-TPT (Yoon et al., 2024), which incorporates calibration-aware objectives. Although effective, these strategies typically require optimization loops, multiple augmentations, and backpropagation through large encoders, resulting in substantial computational and memory overhead at inference. In contrast, VS2 performs a single forward pass with sparse SAE-guided steering, avoiding test-time optimization while still providing measurable gains.
@@ 3 Method
Preliminaries.
A Sparse Autoencoder (SAE) consists of an encoder-decoder pair that maps an input vector $\bm{x}$} into a latent representation $\bm{z}$}, then reconstructs $\bm{x}$} from $\bm{z}$}. Formally, we define the encoder as: ${\bm{z}}=\mathrm{ReLU}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})+{\bm{b}}_{\mathrm{enc}}),$ where ${\bm{W}}_{\mathrm{enc}}\in{\mathbb{R}}^{n\times d}$ is the encoder weight matrix, ${\bm{b}}_{\mathrm{enc}}\in{\mathbb{R}}^{n}$ is a bias term, and ${\bm{b}}_{\mathrm{pre}}\in{\mathbb{R}}^{d}$ normalizes the input. The decoder then reconstructs $\bm{x}$} using a decoder weight matrix , as $\hat{{\bm{x}}}={\bm{W}}_{\mathrm{dec}}\,{\bm{z}}+{\bm{b}}_{\mathrm{pre}}$ . To enforce sparsity, $\ell_{1}$ norm with strength $\alpha$ is introduced as the regularization term, and the intact optimization objective becomes $\mathcal{L}_{\mathrm{SAE}}=\|{\bm{x}}-\hat{{\bm{x}}}\|_{2}^{2}+\alpha\,\|{\bm{z}}\|_{1}.$ In practice, instead of relying on $\ell_{1}$ regularization, top-$k$ Sparse Autoencoder  (Gao et al., 2024a) explicitly enforces the sparsity by using a top-$k$ selection mechanism: ${\bm{z}}=\mathrm{TopK}({\bm{W}}_{\mathrm{enc}}\,({\bm{x}}-{\bm{b}}_{\mathrm{pre}})),$ where $\mathrm{TopK}(\cdot)$ operator retains only the $k$ highest-magnitude activations and zeroes out the rest. This modification ensures that only a fixed number of latent dimensions contribute to the reconstructed representation, making the extracted features more interpretable.
@@ 3.1 Visual Sparse Steering
Visual Sparse Steering (VS2) is a lightweight, label-free, test-time method that guides vision models using steering vectors derived from sparse features learned by top-k Sparse Autoencoders without requiring contrastive data. We hypothesize that the sparse features in an SAE’s latent space can capture salient aspects of the input, thereby reducing redundancy. To this end, VS2 constructs a steering vector by amplifying the SAE’s salient features, reconstructing the corresponding amplified representation, and subtracting the original reconstruction (without amplification) to minimize the inherent SAE’s reconstruction errors. The resulting sparse steering vector points in the direction of the salient features and is linearly combined with the original embedding to steer the model’s behavior. Unlike common steering vectors, VS2 does not require positive and negative anchor examples; instead, steering occurs in the direction of the most salient features identified by the SAE.